A 39 g ice cube at -69°C is placed in a lake whose temperature is 65°C. Calculate the change in entropy of the cube-lake system as the ice cube comes to thermal equilibrium with the lake. The specific heat of ice is 2220 J/kg-K. (Hint: Will the ice cube affect the temperature of the lake?) Numbe UnitšT J/k the tolerance is +/-2% Click if you would like to Show Work for this question: Open Show Work
Answer #1

s = Q/T ... by definition of entropy

Q = mc dT = 0.039 x 2220 x (65-(-69))

= 11601.7 J

T= 65+273 = 338K

s = 34.32 J/K